2. An FM radio station is broadcasting at 106.3 MHz What is the wavelength of the radio wave?
Answer #1

F=106.3Mhz

=106.3*10^6 Hz

f=c/L

c is speed of light

L=wavelength

f=frequecy

L=c/f

=3*10^8/106.3*10^6

=0.0282*10^2

=2.82m